# On the approximability of random-hypergraph MAX-3-XORSAT problems with quantum algorithms

### v1

@@ Abstract
A huge range of important problems in computer science–including task optimization, formal logic, encryption, and machine learning–can be solved by finding the sequence of binary variables that optimizes a cost function defined by a series of few-variable constraint relationships. These problems define the complexity class NP, and are in the worst case, and often the typical case, exponentially hard in the number of variables for all known methods. This hardness applies both to exact and approximate optimization, e.g. finding configurations with a value within a defined fraction of the global optimum. Fundamentally, the lack of any guided local minimum escape method ensures the hardness of both exact and approximate optimization classically, but the intuitive mechanism for approximation hardness in quantum algorithms based on Hamiltonian time evolution is not well understood. In this work, using the prototypically hard MAX-3-XORSAT problem class, we explore this question. We conclude that the mechanisms for quantum exact and approximation hardness are fundamentally distinct. We qualitatively identify why traditional methods such as high depth quantum adiabatic optimization algorithms are not reliably good approximation algorithms. We propose a new spectral folding optimization method that does not suffer from these issues and study it analytically and numerically. We consider random rank-3 hypergraphs including extremal planted solution instances, where the ground state satisfies an anomalously high fraction of constraints compared to truly random problems. We show that, if we define the energy to be $E=N_{\mathrm{unsat}}-N_{\mathrm{sat}}$, then spectrally folded quantum optimization will return states with energy $E\leq AE_{\mathrm{GS}}$ (where $E_{\rm GS}$ is the ground state energy) in polynomial time, where conservatively, $A\simeq 0.6$. We thoroughly benchmark variations of spectrally folded quantum optimization for random classically approximation-hard (planted solution) instances in simulation, and find performance consistent with this prediction. We do not claim that this approximation guarantee holds for all possible hypergraphs, though our algorithm’s mechanism can likely generalize widely. These results suggest that quantum computers are more powerful for approximate optimization than had been previously assumed.
@@ I Introduction
To make further progress, a host of heuristic quantum algorithms have been developed, such as analog quantum annealing Finnila et al. (1994); Kadowaki and Nishimori (1998); Das and Chakrabarti (2008); Johnson et al. (2011); Boixo et al. (2014); Hauke et al. (2019); King et al. (2023), and its closed system and digital cousins, adiabatic quantum computing (AQC) Farhi et al. (2000); Albash and Lidar (2017) and quantum approximate optimization algorithms (QAOAs) Farhi et al. (2014). In all of these methods the spin glass problem Hamiltonian (diagonal in the computational $z$ basis) is combined with a transverse field term (typically, a uniform field along $x$), which allows the system to escape from local minima by multiqubit tunneling, a collective process where large clusters of qubits all change configuration simultaneously to tunnel from one minimum to another Grattan et al. (2023). In some artificial problem instances this process has been shown to produce exponential speedups compared to classical simulated annealing, with polynomial speedups observed in experiment for short-ranged graphs Albash and Lidar (2018); King et al. (2018); Ebadi et al. (2022); King et al. (2023). However, generally applicable beyond-quadratic speedups for NP optimization problems have not been realized, and given the present noisy state of quantum hardware and the projected severe overhead of error correction Babbush et al. (2021), this degree of speedup is insufficient for practical advantage over classical machines.
@@
  • Understanding why this is the case suggests a novel spectral folding quantum strategy, that transforms $H_{P}$ and then solves the transformed Hamiltonian through more traditional means. For some variations the performance of spectrally folded quantum optimization can be predicted analytically and promises an efficient approximation guarantee for an extremely large fraction of instances, well into the classically hard regime. For random hypergraphs in this regime, spectrally folded quantum optimization provides an exponential speedup for returning low energy states.
@@
We back up all of these claims with extensive theoretical analysis and numerical tests of all core predictions, to the largest feasible system sizes for simulating quantum algorithms, and to the largest sizes needed to ensure asymptotic scaling has been reached for classical algorithms. In doing so, we define instance construction rules that ensure classical approximation hardness, at least for algorithms based on local updates, and numerically establish that high depth QAOA does not exhibit meaningful quantum advantage in finding either exact or approximate solutions to these hard instances. We analytically and numerically show that spectrally folded quantum optimization can efficiently approximate these problems, in practice in a linearly growing number of cost function evaluations.
@@
This paper is structured as follows. In section II, we first provide an overview of classical and quantum approximation hardness and the MAX-3-XORSAT problem, and define the construction rules for the problems studied in this work. In section III, we define spectrally folded quantum optimization and propose two variations of our algorithm. Then, in section IV, we develop an extensive-order resummed quantum perturbation theory capable of predicting the performance of these algorithms, and establish an approximation guarantee for random hypergraphs. To verify all of these claims, in section V, we present extensive numerical tests and simulations for a variety of algorithms and problem parametrizations. A summary of the key simulation results is presented in table 1. We finally offer concluding remarks, and include additional technical details in the appendices.
@@ II.1 Inability of previous methods to find or even approach the ground state at large $N$
We emphasize that we are concerned here with an approximation guarantee, not just a method that works well in practice. For completely random problems, one can often predict the average ground state energy using statistical physics arguments, such as the famous Parisi solution to the Sherrington-Kirkpatrick model Parisi (1979). If one is able to find configurations close to this energy in a given instance, that is not sufficient to rule out the existence of some other, much deeper minimum far away in configuration space, even if randomly drawing problems with such deep minima is exponentially unlikely. In other words, there is no efficient classical algorithm to know if a given instance is extremal in this way unless P=NP Arora and Barak (2009).
@@ II.2 Mechanisms for quantum solution hardness: exponentially small gaps and transverse field chaos
These algorithms are fundamentally distinct from classical approaches in two ways. The first is the presumed mechanism for quantum advantage: multiqubit quantum tunneling (or collective quantum phase transitions in general), where local minima can be efficiently escaped through many-body quantum effects that have no classical analog. Second and more subtly, where classical local update algorithms all start from a random high energy state and attempt to cool to low energy states, quantum algorithms start below the energy of the problem ground state. They then attempt to tunnel into the global minimum and into other low energy states as the energy of the initial state crosses the target state as the total Hamiltonian changes.
@@
Unfortunately, when compared to other applications such as quantum simulation or factoring large numbers with Shor’s algorithm, the realistic performance advantage of these algorithms is generally much more modest. In the worst case–and for many problem classes, the typical case–the macroscopic quantum tunneling rate into the ground state decreases exponentially in $N$. This is a fairly generic expectation, as in many cases, including the MAX-3-XORSAT problem discussed below, the tunneling rate at the crossing point can be computed using $N$th order perturbation theory and the convergence of such a method implies exponential decay. For some specific problems this can be circumvented by introducing additional terms into the evolution, such as the well-known result of ramping transverse field terms down one-by-one in mean-field $p$-spin ferromagnets Susa et al. (2018), though generalizations of this and other methods Hauke et al. (2019); Graß (2019); Sels and Polkovnikov (2017); Bao et al. (2018); Yang et al. (2017); Čepaitė et al. (2023) to realistic disordered problems are not expected to show similar advantages at large scales.
@@
Further, for a given class, even if one can somehow ensure that the paramagnet-to-spin-glass transition decays polynomially (as it does in the Sherrington-Kirkpatrick problem Montanari (2021); Farhi et al. (2022) and, likely, for MAXCUT Boulebnane and Montanaro (2021); Basso et al. (2021)), that is not sufficient to ensure that the solution can be found in polynomial time. This is because of a phenomenon known as transverse field chaos (TFC) Altshuler et al. (2010); Knysh (2016), where energetic corrections from the transverse field can change the energy hierarchy of classical minima in the quantum spin glass phase, potentially pushing local minima below the energy of the true ground state of $H_{P}$ (when all transverse terms are turned off). Consequently, optimization methods will steer the system toward these false ground states first, with additional phase transitions that occur as the transverse field is further weakened. As these transitions occur at weak field values, from the analysis in section IV they are generically exponentially slow, and indeed, engineering this effect intentionally is an elegant way to craft hard benchmark problems for quantum algorithms Tang and Kapit (2021). This effect can be avoided by restricting the algorithm to unstructured driver Hamiltonians Farhi et al. (2010) or very weak transverse fields, but in either case performance is very poor. The combination of exponentially small gaps and TFC make it extremely unlikely that any quantum algorithm of this type can reliably and directly find the ground state of NP-complete problems.
@@ II.3 Approximation hardness and conventional AQC/QAOA
Approximation hardness, however, is another story. For many NP-hard problems guaranteeing an approximation better than random guessing by a constant fraction is also NP-hard Sahni and Gonzalez (1976); Håstad (2001); Arora and Barak (2009). As TFC involves crossings between states that were close in energy to begin with, it cannot by itself lead to quantum approximation hardness at this level. So for example, if the random state energy of a given class is chosen to be zero and the ground state $-N$, a sufficiently general convention, then any algorithm which could return states in polynomial time with energy $\leq-cN$ for constant $c>0$ for all instances would promise a potentially exponential speedup. Thus, a hypothetical algorithm which always returned states with energy $-N/3$ or below in polynomial time assuming spectral continuity would still promise an exponential speedup for the hardest problem classes even if TFC reduced that guarantee to $-N/4$. In other words, while TFC can prove ruinous for finding an exact solution when the problem exhibits a clustering phase Mézard et al. (2002, 2005); Hartmann and Weigt (2006); Krzakala and Kurchan (2007); Altarelli et al. (2008); Ibrahimi et al. (2012); Jones et al. (2022) and there are exponentially many states very close to the ground state in energy but well separated from it in Hamming distance, it is not going to push zero energy states into competition with $O\left(1\right)$ fractions of the ground state energy. [^1] It’s important to note that for some problems where the classical approximation hardness threshold is relatively close to 1, such as MAXCUT, TFC might well become a serious obstacle to achieving quantum advantage for approximation.
@@
Exponentially small gaps are a more serious problem, but those too cannot so easily be assumed to ensure approximation hardness. This is because the empirical “difficulty exponents” of phase transitions in low-order CSPs are often quite small; for random hypergraph MAX-3-XORSAT instances, the minimum gap at the paramagnet-spin glass transition scales as approximately $\Omega_{0}\left(N\right)\sim 2^{-cN}$ with $c\simeq 0.14$ (see section IV.2). And while the fraction of states $p_{E}\left(N\right)$ below the approximation threshold is typically exponentially small, the total number of such states is exponentially large. If we could assume that the mixing rate between the paramagnetic initial state and the dressed excited states of $H_{P}$ is equivalent to that of the ground state, then the naive product of $\Omega_{0}^{2}\left(N\right)\times 2^{N}\times p_{E}\left(N\right)$ is often exponentially large as well, which would ensure fast approximation. But this naive analysis is inaccurate, because as we argue momentarily, mixing matrix elements with excited states decay more quickly, and the decay exponents grow with excitation energy. The exponential number of target states in approximation problems allow an algorithm suffering from both TFC and exponentially decaying matrix elements to still guarantee an efficient approximation, provided, at least, that those decay exponents are not too small. In some sense, our novel spectrally folded quantum algorithm achieves fast approximation of classically hard instances by ensuring that the exponential decay rate of tunneling into the problem’s excited states matches or exceeds that of the ground state in more traditional methods.
@@
Given all this, should we expect that AQC or QAOA will efficiently approximate NP-hard problems? Formally, it has not been proven that these algorithms are not efficient approximators when the circuit depth is allowed to grow as a low-order polynomial in system size, though there are good reasons to be doubtful, supported by a number of recent works Basso et al. (2022); Boulebnane and Montanaro (2022); Anshu and Metger (2023); Benchasattabuse et al. (2023). Let us consider, qualitatively, how AQC/QAOA solves such a problem, assuming the limit of quasi-continuous time; we present a quantiative analysis below in Sec. IV.2. The system is initialized in the paramagnetic ground state of a uniform transverse field Hamiltonian $H_{D}$, and the system evolves in time interpolating between $H_{D}$ and $H_{P}$ by lowering the coefficient of one and raising the coefficient of the other, raising the energy of the paramagnetic state until it crosses the problem ground state from below. We again emphasize the fundamental distinction of crossing from below, rather than cooling from above, in quantum and classical algorithms. The minimum gap at the transition is expected to decay exponentially in $N$ and is approximately given by the overlap of the dressed problem ground state $\left|G_{D}\right>$ with the paramagnet state $\left|S\right>$; we perform this calculation in Sec. IV.2. If the dressings are weak enough that $\left|G_{D}\right>=\left|G\right>$ (the Grover limit Roland and Cerf (2002)), then $S|G_{D}\right>=2^{-N/2}$; however for random MAX-3-XORSAT problems the perturbative corrections spread $\left|G_{D}\right>$ over a more significant (if still exponentially small) fraction of Hilbert space and reduce the decay exponent to around a quarter of that in the Grover case.
@@
Assuming that we evolve time too quickly (e.g. not exponentially long) and miss the primary phase transition, we can ask how efficiently $\left|S\right>$ will mix with the problem’s excited states, as these rates ultimately determine the algorithm’s efficacy as an approximator. Calculating them directly is very difficult, but we can qualitatively predict that they should be much smaller for two reasons. First, these crossings occur as the transverse field strength $\kappa$ is reduced toward zero, and since the dressings that reduce the decay exponent all scale with extensive powers of $\kappa$, they will be much reduced by any reduction in $\kappa$ itself. Second, when considering a dressed excited state $\left|E_{D}\right>$, the perturbative dressings that come from mixing with states with lower energy now have opposite sign and destructively interfere with other corrections in the overlap with $\left|S\right>$, in contrast to $\left|G_{D}\right>$ where all higher order terms are positive definite. Both of these effects can, and do, considerably worsen the mixing rates with excited states and make AQC/QAOA a poor approximator in the worst cases, and often in practice.

### v3

@@ Abstract
A canonical feature of the constraint satisfaction problems in NP is approximation hardness, where in the worst case, finding sufficient-quality approximate solutions is exponentially hard for all known methods. Fundamentally, the lack of any guided local minimum escape method ensures both exact and approximate classical approximation hardness, but the equivalent mechanism(s) for quantum algorithms are poorly understood. For algorithms based on Hamiltonian time evolution, we explore this question through the prototypically hard MAX-3-XORSAT problem class. We conclude that the mechanisms for quantum exact and approximation hardness are fundamentally distinct. We review known results from the literature, and identify mechanisms that make conventional quantum methods (such as Adiabatic Quantum Computing) weak approximation algorithms in the worst case. We construct a family of “spectrally filtered” quantum algorithms that escape these issues, and develop analytical theories for their performance. We show that, for random hypergraphs in the approximation-hard regime, if we define the energy to be $E=N_{\mathrm{unsat}}-N_{\mathrm{sat}}$, spectrally filtered quantum optimization will return states with $E\leq q_{m}E_{\mathrm{GS}}$ (where $E_{\rm GS}$ is the ground state energy) in sub-quadratic time, where conservatively, $q_{m}\simeq 0.59$. This is in contrast to $q_{m}\to 0$ for the hardest instances with classical searches. We test all of these claims with extensive numerical simulations. We do not claim that this approximation guarantee holds for all possible hypergraphs, though our algorithm’s mechanism can likely generalize widely. These results suggest that quantum computers are more powerful for approximate optimization than had been previously assumed.
@@ I Introduction
To make further progress, a host of heuristic quantum algorithms have been developed, such as analog quantum annealing Finnila et al. (1994); Kadowaki and Nishimori (1998); Das and Chakrabarti (2008); Johnson et al. (2011); Boixo et al. (2014); Hauke et al. (2019); King et al. (2023), and its closed system and digital cousins, adiabatic quantum computing (AQC) Farhi et al. (2000); Albash and Lidar (2017) and quantum approximate optimization algorithms (QAOAs) Farhi et al. (2014). In all of these methods the spin glass problem Hamiltonian (diagonal in the computational $z$ basis) is combined with a transverse field term (typically, but not exclusively, a uniform field along $x$), which allows the system to escape from local minima through collective quantum fluctuations. In many cases this takes the form of multiqubit tunneling, a collective process where large clusters of qubits all change configuration simultaneously to tunnel from one minimum to another Grattan et al. (2023). In some artificial problem instances these processes have been shown to produce exponential speedups compared to classical simulated annealing, with polynomial speedups observed in experiment for short-ranged graphs Albash and Lidar (2018); King et al. (2018); Ebadi et al. (2022); King et al. (2023); Bauza and Lidar (2024). However, generally applicable beyond-quadratic speedups for NP optimization problems have not been realized, and given the present noisy state of quantum hardware and the projected severe overhead of error correction Babbush et al. (2021), this degree of speedup is insufficient for practical advantage over classical machines.
@@
  • Understanding why this is the case suggests a novel spectrally filtered quantum optimization strategy, that transforms $H_{P}$ through the application of nonlinear filter functions and then solves the transformed Hamiltonian through more traditional means. For some variations the performance of spectrally filtered quantum optimization can be predicted analytically and promises an efficient approximation guarantee for an extremely large fraction of instances, well into the classically hard regime. For random hypergraphs in this regime, spectrally filtered quantum optimization provides an exponential speedup for returning low energy states.
@@
We support these claims with extensive theoretical analysis and numerical tests of all core predictions, to the largest feasible system sizes for simulating quantum algorithms, and to the largest sizes needed to ensure asymptotic scaling has been reached for classical algorithms. In doing so, we define instance construction rules that ensure classical approximation hardness, at least for algorithms based on local updates, and numerically establish that high depth Trotterized AQC (TAQC) does not exhibit meaningful quantum advantage in finding either exact or approximate solutions to these hard instances. We analytically and numerically show that spectrally filtered quantum optimization can efficiently approximate these problems, in practice in a linearly or sub-quadrically growing number of cost function evaluations, depending on formulation. We present all these results as constructive evidence in support of our core conjecture:
Core conjecture: Quantum approximability of random MAX-3-XORSAT instances – Let $\mathcal{P}\left(N_{C},N,\epsilon\right)$ be the set of all MAX-3-XORSAT instances with $N_{C}$ unique three-body constraints (with no negations), $N$ binary variables, and where the optimal configurations $G$ satisfy $\left(1-\epsilon\right)N_{C}$ constraints in total. Let $E\left(s\right)=-N_{C}\left(n_{\mathrm{sat}}\left(s\right)-n_{\mathrm{unsat}}\left(s\right)\right)$ be our definition of the classical energy of configuration $s$. Let $\epsilon\ll 1$ be a small constant, let the density of constraints $N_{C}/N\equiv d_{C}\gg 1$ be large, and let $q_{m}>0$ be a constant independent of $\epsilon$ and $d_{C}$. Finally, let $H_{P}$ be a random instance drawn from $\mathcal{P}\left(N_{C},N,\epsilon\right)$. Then, with high probability, but not guaranteed for all instances, a quantum computer with $O\left(N\right)$ noise-free qubits can return string(s) with $E\leq q_{m}E\left(G\right)$ in $O\left(N_{C}^{2}N^{2}{\rm polylog}\left(N\right)\right)$ or fewer total gates. We further conjecture that $q_{m}\geq 0.6$, and that for $\epsilon\ll 1$ and $d_{C}\gg 1$ it is likely close to 1.
We present two algorithms which would prove this conjecture true if the analytically predicted scaling of either one holds asymptotically as $N\to\infty$. The algorithms are both based on spectral filtering, but their structure, and the methods used to predict performance, are very different; we see the fact that two very different calculations based on different methods return similar results as further suggestive that our conjecture is true.
The central observation underlying both approaches is that if one is focused on finding approximate solutions using quantum time evolution, the difficulty is controlled mostly by global statistical properties of the problem, many of which are hypergraph-independent. For a “direct” implementation of the MAX-3-XORSAT cost function, these properties imply an exponential time to solution. However, by optimizing a filtered cost function, $f\left(H_{P}\right)$ (where $f\left(E\right)$ is any real scalar function of the classical energy), one can tune the most important statistical properties and there are straightforward choices of $f\left(E\right)$ that lead to low-order polynomial scaling, at least for random hypergraphs. We state the above claim as a conjecture rather than a proved statement owing to the fact that both analytical calculations rely on approximation steps to make them tractable, and while these approximations are well-supported theoretically and the final results are matched by numerics up to the largest system sizes we could computationally access, we err on the side of caution nonetheless.
@@
This paper is structured as follows. In section II, we first provide an overview of classical and quantum approximation hardness and the MAX-3-XORSAT problem, define the construction rules for the problems studied in this work, and state the principles which guide the parameter choices we make in formulating our algorithms. In section III, we define spectrally filtered quantum optimization and propose three optimization algorithms based on it. Then, in section IV, we develop novel theoretical tools capable of predicting the performance of these algorithms, and establish an approximation guarantee for random hypergraphs. To verify all of these claims, in section V, we present extensive numerical tests and simulations for a variety of algorithms and problem parametrizations. A summary of the key simulation results is presented in table 1. We finally offer concluding remarks, and include additional technical details in the appendices.
@@ II.1 Inability of previous methods to find or even approach the ground state at large $N$
We emphasize that we are concerned here with an approximation guarantee, not just a method that works well in practice. For completely random problems, one can often predict the average ground state energy using statistical physics arguments, such as the famous Parisi solution to the Sherrington-Kirkpatrick model Parisi (1979). If one is able to find configurations close to this energy in a given instance, that is not sufficient to rule out the existence of some other, much deeper minimum far away in configuration space, even if randomly drawing problems with such deep minima is exponentially unlikely. In other words, there is no efficient classical algorithm to know if a given instance is extremal in this way unless P=NP Arora and Barak (2009) (see section II.4 for more details).
@@ II.2 Mechanisms for quantum solution hardness: exponentially small gaps and transverse field chaos
These algorithms are fundamentally distinct from classical approaches in two ways. The first is the presumed mechanism for quantum advantage: collective quantum phase transitions (including multiqubit tunneling), where local minima can be efficiently escaped through many-body quantum effects that have no classical analog. Second and more subtly, where classical local update algorithms all start from a random high energy state and attempt to cool to low energy states, the quantum algorithms considered in this work start below the energy of the problem ground state. They then attempt to transition into the global minimum and into other low energy states as the energy of the initial state crosses the target state from below as the total Hamiltonian changes.
@@
Unfortunately, when compared to other applications such as quantum simulation or factoring large numbers with Shor’s algorithm, the realistic performance advantage of these algorithms is generally much more modest. In the worst case–and for many problem classes, the typical case–the macroscopic quantum tunneling rate into the ground state decreases exponentially in $N$. This is a fairly generic expectation, as in many cases, including the MAX-3-XORSAT problem discussed below, the tunneling rate at the crossing point can be computed using $N$th order perturbation theory and the convergence of such a method implies exponential decay.
@@
But let us imagine that one could somehow ensure that the gap at the paramagnet-to-spin-glass transition decays polynomially in size of the problem. This occurs naturally in the Sherrington-Kirkpatrick problem Montanari (2021); Farhi et al. (2022) and, likely, for MAXCUT Boulebnane and Montanaro (2021); Basso et al. (2021), and in other classes it can sometimes be engineered through clever algorithm formulation Susa et al. (2018); Hauke et al. (2019); Graß (2019); Sels and Polkovnikov (2017); Bao et al. (2018); Yang et al. (2017); Čepaitė et al. (2023). Unfortunately, even this is not sufficient to ensure that the solution can be found in polynomial time. This is because of a phenomenon known as transverse field chaos (TFC) Altshuler et al. (2010); Knysh (2016), where energetic corrections from the transverse field can change the energy hierarchy of classical minima in the quantum spin glass phase, potentially pushing local minima below the energy of the true ground state of $H_{P}$ (when all transverse terms are turned off). Consequently, optimization methods will steer the system toward these false ground states first, with additional phase transitions that occur as the transverse field is further weakened. As these transitions occur at weak field values, from the analysis in section IV they are generically exponentially slow, and indeed, engineering this effect intentionally is an elegant way to craft hard benchmark problems for quantum algorithms Tang and Kapit (2021). This effect can be avoided by restricting the algorithm to unstructured driver Hamiltonians Farhi et al. (2010) or very weak transverse fields, but in either case performance is very poor. The combination of exponentially small gaps and TFC make it extremely unlikely that any quantum algorithm of this type can reliably and directly find the ground state of NP-complete problems.
@@ II.3 Approximation hardness and conventional AQC/QAOA
Approximation hardness, however, is another story. For many NP-hard problems guaranteeing an approximation better than random guessing by a constant fraction is also NP-hard Sahni and Gonzalez (1976); Håstad (2001); Arora and Barak (2009). As TFC involves crossings between states that were close in energy to begin with, it cannot by itself lead to quantum approximation hardness at this level. So for example, if the random state energy of a given class is chosen to be zero and the ground state $-N$, a sufficiently general convention, then any algorithm which could return states in polynomial time with energy $\leq-cN$ for constant $c>0$ for all instances would promise a potentially exponential speedup. Thus, a hypothetical algorithm which always returned states with energy $-N/3$ or below in polynomial time assuming spectral continuity would still promise an exponential speedup for the hardest problem classes even if TFC reduced that guarantee to $-N/4$. In other words, while TFC can prove ruinous for finding an exact solution when the problem exhibits a clustering phase Xu et al. (2020); Mézard et al. (2002, 2005); Hartmann and Weigt (2006); Krzakala and Kurchan (2007); Altarelli et al. (2008); Ibrahimi et al. (2012); Jones et al. (2022) and there are exponentially many states very close to the ground state in energy but well separated from it in Hamming distance, it is not going to push zero energy states into competition with $O\left(1\right)$ fractions of the ground state energy. [^1] It’s important to note that for some problems where the classical approximation hardness threshold is relatively close to 1, such as MAXCUT, TFC might well become a serious obstacle to achieving quantum advantage for approximation.
@@
Exponentially small gaps are a more serious problem, but those too cannot so easily be assumed to ensure approximation hardness. This is because the empirical “difficulty exponents” of phase transitions in low-order CSPs are often quite small; for random hypergraph MAX-3-XORSAT instances, the minimum gap at the paramagnet-spin glass transition scales as approximately $\Omega_{0}\left(N\right)\sim 2^{-cN}$ with $c\simeq 0.125-0.14$ (see section IV.2). Quite generally (see the MSCALE conjecture in Ref. Kapit and Oganesyan (2021)) this is an asymptotically accurate proxy for the tunneling matrix element, which we can use to approximate the tunneling rate from the paramagnetic initial state and the dressed excited states of $H_{P}$. Crucially, while the fraction of states $p_{E}\left(N\right)$ below the approximation threshold is typically exponentially small, the total number of such states is exponentially large, $\sim 2^{N}\times p_{E}\left(N\right)$. Thus an optimistic first approximation would be to look for an energy $E$ where $\Omega_{0}^{2}\left(N\right)\times 2^{N}\times p_{E}\left(N\right)$ stops decaying exponentially in $N$, which should be a relatively easy task.
But this naive analysis is inaccurate because, as we explain momentarily, destructive quantum interference and a weakening transverse field suppress matrix elements to excited states as compared to ground state mixing, and the decay exponents grow with excitation energy. The exponential number of target states in approximation problems should still allow an algorithm suffering from both TFC and exponentially decaying matrix elements to guarantee an efficient approximation, provided, at least, that those decay exponents are not too large.  [^2] In a technical sense, the success of the novel spectrally folded trial minimum annealing algorithm developed in this work pivots on the mechanism to suppress destructive quantum interference and maintain a strong transverse field at the transition point, thus preserving or even improving small tunneling rate exponent in the estimate above. It is worth noting that a dimensionless Thouless criterion similar (but distinct) to the rate-based one above has been advocated as an accurate estimate for the many-body localization-delocalization transitionPal and Huse (2010); Serbyn et al. (2017) – localization requires the matrix element for tunneling to be smaller than the level spacing for which inverse density of states is often an accurate proxy. While we generally expect excited states above a typical local minumum to be (intra-well) delocalized, they are not inter-well delocalized and so the question of efficient approximability we are interested in is likely closely related to whether or not we are able to delocalize the many-body dynamics across different local minima, particularly for the trial minimum annealing formulation in section III.4. Note, however, that of the two criteria the rate-based one is the more conservative of the two. We intend to revisit this analogy between approximability and delocalization in the future.
@@
Given all this, should we expect that AQC or QAOA will efficiently approximate NP-hard problems? Formally, it has not been proven that these algorithms are not efficient approximators when the circuit depth is allowed to grow as a low-order polynomial in system size, though there are good reasons to be doubtful, supported by a number of recent works Basso et al. (2022); Boulebnane and Montanaro (2022); Anshu and Metger (2023); Benchasattabuse et al. (2023). Let us consider, qualitatively, how AQC and related methods solve such a problem, assuming the limit of quasi-continuous time; we present a quantiative analysis below in Sec. IV.2. The system is initialized in the paramagnetic ground state of a uniform transverse field Hamiltonian $H_{D}$, and the system evolves in time interpolating between $H_{D}$ and $H_{P}$ by lowering the coefficient of one and raising the coefficient of the other, raising the energy of the paramagnetic state until it crosses the problem ground state from below. We again emphasize the fundamental distinction of crossing from below, rather than cooling from above, in quantum and classical algorithms here. The minimum gap at the transition is expected to decay exponentially in $N$ and is approximately given by the overlap of the dressed problem ground state $\left|G_{D}\right>$ with the paramagnet state $\left|S\right>$; we perform this calculation in Sec. IV.2. If the dressings are weak enough that $\left|G_{D}\right>=\left|G\right>$ (the Grover limit Roland and Cerf (2002)), then $S|G_{D}\right>=2^{-N/2}$; however for random MAX-3-XORSAT problems the perturbative corrections spread $\left|G_{D}\right>$ over a more significant (if still exponentially small) fraction of Hilbert space and reduce the decay exponent to around a quarter of that in the Grover case.
@@
Assuming that we evolve time too quickly (e.g. not exponentially long) and miss the primary phase transition, we can ask how efficiently $\left|S\right>$ will mix with the problem’s excited states, as these rates ultimately determine the algorithm’s efficacy as an approximator. Calculating them directly is very difficult, but we can qualitatively predict that they should be much smaller for two reasons. First, these crossings occur as the transverse field strength $\kappa$ is reduced toward zero, and since the dressings that reduce the decay exponent all scale with extensive powers of $\kappa$, they will be much reduced by any reduction in $\kappa$ itself. Second, when considering a dressed excited state $\left|E_{D}\right>$, the perturbative dressings that come from mixing with states with lower energy now have opposite sign and destructively interfere with other corrections in the overlap with $\left|S\right>$, in contrast to $\left|G_{D}\right>$ where all higher order terms are positive definite. Both of these effects can, and do, considerably worsen the mixing rates with excited states and make AQC a poor approximator in the worst cases, and often in practice. That said, we do note a recent encouraging result demonstrating an empirical scaling advantage over classical heuristics in finding approximate solutions to 2d spin glasses, using real quantum annealing hardware Bauza and Lidar (2024).